1. 2. Suppose the below reaction in the forward direction is first order in A and the rate constant is 1.10×10-2 s–1. A ⇌ B. The reverse reaction is first order in B and the rate constant is 4.70×10-2 s–1 at the same temperature.
What is the value of the equilibrium constant for the below reaction at this temperature?
A ⇌ B
1
Expert's answer
2020-04-02T09:03:06-0400


k1 is rate constant of forward direction, k1 = 1.10*10-2 s-1

k-1 is rate constant of backward direction, k-1 = 4.70*10-2 s-1

equilibrium constant K = "{\frac {k_1} {k_{-1}} }" = "{\frac {1.10*10^{-2}} {4.70*10^{-2}}}" =0.234 = 2.34*10-1